OpenStudy (anonymous):

Use differentials to approximate (3.98)^4. How?

OpenStudy (amistre64):

well, a linear approximation might help

OpenStudy (anonymous):

Here's an idea... start with \[\huge y = f(x) = x^4\]

OpenStudy (amistre64):

say y = (x+c)^4 y' = 4(x+c)^3

OpenStudy (amistre64):

y = y'(x+c) - y(c) is a linear approximation

OpenStudy (anonymous):

\[\huge \frac{dy}{dx}=4x^3\]\[\huge dy = 4x^3 dx\]